OpenStudy (anonymous):

The system of equations represented by x+2y=2 & -4x+3y=25 intersects at (a,b). What does -2(a+b)=?

jimthompson5910 (jim_thompson5910):

what did you get for a and b

jimthompson5910 (jim_thompson5910):

unfortunately that is incorrect

jimthompson5910 (jim_thompson5910):

sorry one sec

jimthompson5910 (jim_thompson5910):

first you solve x+2y=2 for x x+2y=2 x=2-2y

jimthompson5910 (jim_thompson5910):

then you plug it into -4x+3y=25 and solve for y -4x+3y=25 -4(2-2y)+3y=25 -8+8y+3y=25 .... .... .... y = ???

jimthompson5910 (jim_thompson5910):

correct, now use that to find x

jimthompson5910 (jim_thompson5910):

perfect

jimthompson5910 (jim_thompson5910):

you got it

jimthompson5910 (jim_thompson5910):

so a = -4, b = 3

jimthompson5910 (jim_thompson5910):

very nice

jimthompson5910 (jim_thompson5910):

you're welcome